# TwinBrainVLA[Uncaptioned image]: Unleashing the Potential of Generalist VLMs for Embodied Tasks via Asymmetric Mixture-of-Transformers

### v1

@@ 5.2 Experiments on RoboCasa
[…]
@@
[…]
Baselines. We compare TwinBrainVLA against several competitive baselines trained on the same dataset. Specifically, we include Isaac-GR00T-N1.6 (Team et al., 2025b), a large-scale VLA model that demonstrates strong performance on humanoid manipulation tasks. Additionally, to directly assess the advantages of our dual-brain architecture, we benchmark against two representative models from the starVLA framework (starVLA, 2025): QwenGR00T and QwenPI, both instantiated with Qwen3-VL-4B-Instruct backbones. These baselines follow similar training protocols and utilize the same dataset, enabling a fair comparison and highlighting the specific contributions of our asymmetric dual-VLM design.
6 Conclusion
In this paper, we introduced TwinBrainVLA[Uncaptioned image], a novel VLA framework that addresses the fundamental tension between semantic generalization and embodied specialization in robotic learning. By drawing inspiration from hemispheric lateralization, we proposed an asymmetric dual-stream architecture that structurally decouples the visual cognitive process. Our AsyMoT mechanism effectively synergizes a frozen "Left Brain" for robust open-world reasoning with a trainable "Right Brain" for precise sensorimotor control, enabling the system to master complex manipulation skills without "catastrophic forgetting" of its general visual capabilities.
Limitation and Future Work
While TwinBrainVLA provides a powerful framework for decoupling semantic and embodied processing, several limitations remain that outline our future research directions.
A more decoupled model architecture. First, the current implementation of the Asymmetric Mixture-of-Transformers (AsyMoT) mechanism requires the Left and Right Brains to share an identical model architecture. This structural constraint ensures compatible hidden state dimensions for seamless joint attention but limits the flexibility to pair heterogeneous models—for instance, synergizing a massive, reasoning-heavy VLM with a lightweight, high-frequency control model. We are actively exploring generalized fusion mechanisms, such as learnable projection layers or cross-attention adapters, to support backbones of varying sizes and architectures.
More diverse model checkpoints. Second, both streams in our current setup are initialized from standard general-purpose VLM checkpoints (e.g., Qwen-VL). However, the field has arguably seen a surge in specialized Embodied VLMs trained specifically for robotics. Integrating these specialized models into our dual-brain paradigm—potentially initializing the "Right Brain" with an embodied checkpoint while retaining a generalist "Left Brain"—presents a promising avenue. We are currently investigating strategies to effectively adapt our framework to leverage these domain-specific post-trained weights.
Training on larger-scale data. Third, regarding data scale, our current training utilizes subsets of the Open X-Embodiment dataset. We acknowledge that the capacity of our dual-stream architecture may not be fully realized with limited data. We are in the process of scaling our pre-training to the complete dataset to further enhance the model’s robustness and downstream performance.
More benchmarks and real-robot experiment. Finally, future work will also focus on extending our evaluation to a broader range of benchmarks, including real-robot scenarios, to comprehensively assess the versatility of TwinBrainVLA.
Appendix A Training Hyperparameters
Appendix B VLM Prompt Template

### v2

@@ Appendix D RoboCasa Tabletop Benchmark
Task Isaac-GR00T N1.6 QwenGR00T + Qwen3VL QwenPI + Qwen3VL TwinBrainVLA + Qwen2.5VL TwinBrainVLA + Qwen3VL
@@
Average 47.6 47.8 43.9 53.5 54.6
Appendix E VLM Prompt Template
@@
Appendix F Implementation Details on the LIBERO Benchmark
Training Dataset To evaluate the model’s multi-task generalization capabilities, we adopted a unified training strategy. We aggregated the training datasets from all four LIBERO task suites (LIBERO-Spatial, -Object, -Goal, and -Long) into a single combined dataset. Instead of training separate experts for each suite, we trained a single generalist model on this unified data.
Training Configuration. We utilized Qwen3-VL-4B as the visual-language backbone. The model was trained on a cluster of 8 NVIDIA H100 GPUs. We set the per-device batch size to 16, resulting in a global batch size of 128. The training process was conducted for a total of 100,000 steps.
Evaluation Protocol. For evaluation, we strictly followed the official LIBERO evaluation protocol. We utilized the standard evaluation scripts provided by the benchmark. The performance was assessed across a total of 500 trials, consisting of 10 distinct tasks with 50 evaluation episodes per task. The final reported result represents the average success rate across these 500 trials.
Appendix G "Twin-to-One" via Knowledge Distillation
G.1 Motivation and Methodology
While TwinBrainVLA achieves superior performance by leveraging a dual-stream architecture to utilize general capabilities, the inclusion of a second VLM inevitably increases computational overhead during deployment. To bridge the gap between high-level capability utilization and deployment efficiency, we propose a "Twin-to-One" knowledge distillation framework.
Our core hypothesis is that the "Right Brain" of TwinBrainVLA, having dynamically fused semantic knowledge from the frozen "Left Brain" via AsyMoT, possesses a superior feature space compared to a standard single-stream VLA. Therefore, we treat the fully trained TwinBrainVLA as the Teacher and a standard Vanilla VLA as the Student. The goal is to force the Student to internalize the Teacher’s fused general capabilities by mimicking its internal representations.
Distillation Objective. We utilize the last hidden states of the Teacher’s Right Brain as the supervision signal. Formally, let $H_{teacher}\in\mathbb{R}^{L\times D}$ denote the final hidden states of the Right Brain in TwinBrainVLA, and $H_{student}\in\mathbb{R}^{L\times D}$ denote the final hidden states of the Student Vanilla VLA. In addition to the standard action prediction loss $\mathcal{L}_{action}$, we introduce a feature alignment loss $\mathcal{L}_{feat}$:
$\mathcal{L}_{feat}=\frac{1}{L}\sum_{i=1}^{L}\|H_{student}^{(i)}-\text{sg}(H_{teacher}^{(i)})\|^{2}_{2}$ (9)
where $\text{sg}(\cdot)$ indicates the stop-gradient operation, ensuring the Teacher remains fixed. The total training objective for the Student is:
$\mathcal{L}_{total}=\mathcal{L}_{action}+\lambda\mathcal{L}_{feat}$ (10)
where $\lambda$ is a balancing coefficient (set to 1.0 in our experiments).
G.2 Experiment Setup
We conducted the distillation experiment on the SimplerEnv benchmark. The training data (Bridge-V2 and Fractal subsets) and hyperparameters (e.g., learning rate, batch size, optimizer) are kept strictly identical to the main experiments described in Appendix B. The Student model architecture is identical to the Right Brain in TwinBrainVLA.
G.3 Results and Analysis
The quantitative results are presented in Table 9.
Table 9: "Twin-to-One" Distillation Results on SimplerEnv. We compare the distilled single-stream model (Student) against the standard Vanilla VLA and the Teacher (TwinBrainVLA). All models use the Qwen3-VL-4B-Instruct backbone.
Method Spoon on Towel Carrot on Plate Stack Block Eggplant in Basket Average
Baselines
Vanilla VLA (Standard Training) 87.5 50.0 29.2 54.2 55.2
Teacher
TwinBrainVLA (dual-stream) 87.5 58.3 33.3 79.1 64.5
"Twin-to-One"
Distilled Student (single-stream) 83.3 54.2 29.2 66.7 58.4
Performance Gains. As shown in Table 9, the Distilled Student achieves an average success rate of 58.4%, outperforming the standard Vanilla VLA (55.2%) by +3.2%.
Conclusion. These results substantiate that the performance gains of TwinBrainVLA stem from the high-quality, semantics-rich features fused from the Left Brain. Furthermore, it demonstrates that TwinBrainVLA can serve as an effective supervisor, enabling single-stream models to internalize general capabilities and break the ceiling of standard fine-tuning paradigms.
Appendix H Limitation and Future Work
Efficiency. The addition of a second VLM component in TwinBrainVLA leads to higher training and deployment costs compared to vanilla VLAs. While our approach prioritizes generalizability over efficiency, we acknowledge these overheads as a limitation. Nevertheless, it is encouraging that our architecture, consisting of two 4B models, outperforms a larger 8B-parameter model, demonstrating the effectiveness of this design.
Towards Deeper Semantic Fusion. Although our architecture successfully enables the retention and retrieval of general capabilities, how to seamlessly fuse this semantic information into downstream VLA tasks for true generalization is not yet fully resolved. This paper identifies a viable pathway towards this goal, but achieving a definitive solution requires further exploration. We believe that more advanced mechanisms for integrating high-level semantics with low-level policies will be a critical area for future investigation.